# Discrete Time Crystal Phase as a Resource for Quantum Enhanced Sensing

### v3

@@ 1 Introduction
2 Quantum parameter estimation
@@ 3 The model
Here $J$ is the spin exchange coupling, and $\sigma^{x,y,z}_{j}$ are the Pauli operators. The gradient $zz$ interaction in $H_{I}$ causes off-resonant energy splitting at each site and, therefore, leads the particle’s wave function to localize, reminiscent of the localization which is usually induced by applying a gradient magnetic field [95, 96, 89, 90, 91]. This localization, characterized by the existence of an extensive number of conserved quantities [97, 98, 99], is essential to prevent our system from absorbing the energy of the periodic drives [100, 101]. In the absence of the localization, any local physical observable becomes featureless, and the system thermalizes [102, 103]. Since $H_{P}$ acts in period $T$, the Floquet unitary operator for one period evolution is
@@
here $\Omega{=}JT$, and $\Phi$ is tuned to be $\Phi{=}(1{-}\varepsilon)\frac{\pi}{2}$, with $\varepsilon$ as deviation from a $\pi/2$ $x$-rotation. In the following, we show how two main parameters, namely $\Omega$ and $\varepsilon$ play roles in establishing a stable DTC. First, we analytically show that setting $\Omega{=}\pi/2$ results in a stable period doubling DTC that is robust against arbitrary imperfection $\varepsilon$. Then, through comprehensive numerical simulations, we show that as $\Omega$ varies from $\pi/2$, the system goes through a sharp phase transition from a stable DTC to a regime in which DTC order is lost. We explore the possibility of this phase transition to act as resource for quantum sensing.
@@
4 Discrete Time Crystal
@@ 7 Non-DTC region
8 Experimental realization
We propose superconducting quantum simulators, similar to Ref. [35], for realizing our DTC. Rewriting the Floquet operator $U_{F}(\Omega,\varepsilon)=e^{-iH_{P}}e^{-i\Omega H_{I}}$ as
@@ 8 Experimental realization
9 Conclusion
We have established a DTC with indefinite persistent oscillations and strong robustness against imperfections in the driving pulse. We show that this DTC has an application for measuring the coupling strength with quantum-enhanced precision over a region that extends over the entire DTC phase. Through finite-size scaling analysis, we characterize the nature of the phase transition as the second-order and determine relevant critical exponents. In addition, we show that imperfection in the pulse, enhances the precision before melting the DTC. The proposed mechanism is independent of the initial state and can be realized in existing quantum simulators. This provides novel application of DTC phases in the field of quantum metrology.
10 Acknowledgement

### v4

@@ I Introduction
II Quantum parameter estimation
@@ III The model
Here $J$ is the spin exchange coupling, and $\sigma^{x,y,z}_{j}$ are the Pauli operators. The gradient $zz$ interaction in $H_{I}$ causes off-resonant energy splitting at each site and, therefore, leads the particle’s wave function to localize, reminiscent of the localization which is usually induced by applying a gradient magnetic field [95, 96, 89, 90, 91]. This localization, characterized by the existence of an extensive number of conserved quantities [97, 98, 99], is essential to prevent our system from absorbing the energy of the periodic drives [100, 101]. In the absence of the localization, any local physical observable becomes featureless, and the system thermalizes [102, 103]. Since $H_{P}$ acts in period $T$, the Floquet unitary operator for one-period evolution is
@@
here $\Omega{=}JT$, and $\Phi$ is tuned to be $\Phi{=}(1{-}\varepsilon)\frac{\pi}{2}$, with $\varepsilon$ as deviation from a $\pi/2$ $x$-rotation. In the following, we show how two main parameters, namely $\Omega$ and $\varepsilon$ play roles in establishing a stable DTC. First, we analytically show that setting $\Omega{=}\pi/2$ results in a stable period doubling DTC that is robust against arbitrary imperfection $\varepsilon$. Then, through comprehensive numerical simulations, we show that as $\Omega$ varies from $\pi/2$, the system goes through a sharp phase transition from a stable DTC to a regime in which DTC order is lost. We explore the possibility of this phase transition to act as a resource for quantum sensing.
@@
IV Discrete Time Crystal
@@ VII Non-DTC region
VIII Experimental realization
To provide a proof-of-principle demonstration of our DTC sensor, we propose superconducting quantum simulators, similar to Ref. [35], for realizing our DTC. Rewriting the Floquet operator $U_{F}(\Omega,\varepsilon)=e^{-iH_{P}}e^{-i\Omega H_{I}}$ as
@@ VIII Experimental realization
IX Effect of Disordered Interaction
So far, we have concentrated on a uniform gradient interaction among the system’s components. In this section, we aim to analyze the effect of disorder in the gradient interaction. We replaced $H_{I}$ in Eq. (1) with $H_{I}{=}\sum_{j=1}^{L-1}j(1{-}d_{j})\sigma^{z}_{j}\sigma^{z}_{j+1}$, where $d_{j}$ is drawn from a uniform distribution $[-D,D]$. The obtained results are presented in Fig. 6 for two values of the disorder strength namely, $D{=}0.01$ and $D{=}0.05$. As can be seen from Fig. 6 (a) and (b) a slight disorder in the gradient interaction can not affect the dynamical behavior of the DTC and, hence, its performance as a sensor. Increasing the disorder strength to $D{=}0.05$ slightly broadens the transition area, see Fig. 6 (c) and (d). Instead of a sharp transition from the DTC phase to the non-DTC region, some nontrivial oscillations blur the phase boundary. These results are obtained for a chain of size $L{=}12$ with $\varepsilon{=}0.01$ and have been averaged over $20$ repetition. This analysis shows that our DTC probe can tolerate a reasonable amount of disorder in the gradient interaction.
Figure 6: (a) Dynamical behavior of the revival fidelity and (b) the QFI at stroboscopic times as a function of $\omega$ obtained for a disordered chain with strength $D{=}0.01$. (c) The revival fidelity and (d) the QFI obtained for a disordered chain while the disorder strength is $D{=}0.05$. In all the panels the size of the chain is $L{=}12$ and we set $\varepsilon{=}0.01$.
X Conclusion
We have established a DTC with indefinite persistent oscillations and strong robustness against imperfections in the driving pulse. We show that this DTC can be used for measuring the coupling strength with quantum-enhanced precision over a region that extends over the entire DTC phase. Through finite-size scaling analysis, we characterize the nature of the phase transition as the second-order and determine relevant critical exponents. In addition, we show that imperfection in the pulse, enhances the precision before melting the DTC. The proposed mechanism is independent of the initial state and can be realized in existing quantum simulators. It also shows robustness to slight disorder in the gradient interaction.
XI Acknowledgement